OpenStudy (anonymous):

solving systems of equations by grafting -x+3y=9

OpenStudy (anonymous):

is there another equation to graph or just that one? the fastest way to graph this is to let x=0, then solve for y. and then let y=0 and solve for x. you'll have two points and you can then draw a line through them. ^_^

OpenStudy (anonymous):

-2x+by=18

OpenStudy (anonymous):

ok great, have you graphed the first one yet like i outlined?

OpenStudy (anonymous):

Sorry, I am trying to help my Son with homework and we are both getting stressed out. The problem is -x+3y=9 -2X+6Y=18

OpenStudy (anonymous):

no worries

OpenStudy (anonymous):

so we have these two equations: -x+3y=9 -2X+6Y=18 we need to get two points for each of these equations. I'll work out the first one so you see what to do: -x+3y=9, let x=0 -(0)+3y=9 3y=9 y=3 first point will be (0, 3) -x+3y=9, let y=0 -x+3(0)=9 -x=9 or x=-9 second point will be (-9,0) then I will put these two points on the graph: 

OpenStudy (anonymous):

are you able to do the second equation with the methods I did above?

OpenStudy (anonymous):

His answer is y=2/6x-3

OpenStudy (anonymous):

correct, except its +3 not the -3 -2X+6Y=18 6Y = +2X + 18 Y = (2/6)X + 3 so with this equation we need to graph it. there are two ways to graph it, I showed you one way above using points. the other way is to use the y-intercept of +3 and the slope of (2/6). you can do either way. we just need to plot the equation onto the graph.

OpenStudy (anonymous):

ok we have this. thank you so much for the help. God bless

OpenStudy (anonymous):

you are welcome